OpenStudy (123ab456c):

A store is selling 5 notebooks for $15.00. At this rate, what is the cost of 8 notebooks? A. $15.05 B. $15.08 C. $15.13 D. $24.00

Mehek (mehek14):

first find how much PER book 15/5 = 3 so $3 per book 3 * 8 = ?

OpenStudy (blm1):

15/5=price each

OpenStudy (blm1):

15/5=? | 8x?=answer Tag me if you need help ;)

OpenStudy (123ab456c):

@blm1 is the answer D

OpenStudy (blm1):

Yep, great job!

OpenStudy (123ab456c):

thanks! i will fan and medal